# Effects of bumblebee gravity on neutrino motion

### v1

@@ VI Numerical investigation
According to the results presented in Refs. neu53 ; Shi:2024flw , the curvature–induced deviation in a neutrino’s path—quantified by the deflection angle $\delta$—is intrinsically linked to the impact parameter $b$. This relationship is expressed through the following equation 14 :
@@
$=b^{2}\left(\dfrac{x_{D}}{r_{S}}+1-\dfrac{\pi\ell y_{D}}{2r_{S}}\right)-\dfrac{4Mb_{y}D}{r_{S}}.$ (56)
We analyzed how neutrino oscillations are affected by the spacetime of a Bumblebee black hole, focusing on gravitational corrections introduced through the Lorentz–violating parameter $\ell$, as we mentioned earlier. Oscillation patterns were compared between the Bumblebee scenario ($\ell\neq 0$) and the Schwarzschild case ($\ell=0$), with Figs. 4, 5, and 6 highlighting how the azimuthal angle $\varphi$ and the mass–squared difference $\Delta m^{2}$ influence flavor transition probabilities.
@@
The influence of $\ell$ is detailed in Fig. 5. Increasing $\ell$ from $1\times 10^{-10}$ to $3\times 10^{-10}$ slightly changes the oscillation frequency but noticeably reduces the whole probability of conversion. This aspect results from a reduction in the effective phase due to radial stretching of the spacetime. It is worthy to be highlight that, when the Kalb–Ramond fields takes into account instead, it did not produce such radial rescaling, showing no comparable suppression Shi:2025rfq .
On the other hand, Fig. 6 addresses the role of the absolute mass scale. As the lightest neutrino mass $m_{1}$ increases from 00 to $0.02\,\mathrm{eV}$, the amplitude decreases for the normal hierarchy and increases for the inverted one—behavior also seen in Kalb–Ramond gravity. However, the two models differ at $\varphi=0$ due to distinct modifications of the metric: Bumblebee spacetime preserves $\mathfrak{g}_{tt}$, whereas Kalb–Ramond modifies it via a 1/($1/(1-\ell)$ factor, changing local energy measurements and initial oscillation conditions.
@@ VII Conclusion
Propagation effects were governed by a reduced oscillation phase $\tilde{\Phi}_{k}=\frac{m_{k}^{2}}{2E_{0}}\left(1-\frac{\ell}{2}\right)(r_{D}+r_{S})\left(1-\frac{b^{2}}{2r_{D}r_{S}}+\frac{2M}{r_{D}+r_{S}}\right)$, so higher values of $\ell$ lengthened the oscillation scale along both radial and lensed paths. Numerical simulations revealed that raising $\ell$ from $1\times 10^{-10}$ to $3\times 10^{-10}$ left the azimuthal period almost unchanged but suppressed the $\nu_{e}\to\nu_{\mu}$ transition amplitude by up to twenty percent, an effect traced to the radial stretching of the optical metric. More so, the inverted mass ordering consistently produced higher conversion probabilities than the normal one with changing the lightest mass from 00 to $0.02\,\mathrm{eV}$.

### v3

@@ VI Numerical investigation
According to the results presented in Refs. neu53 ; Shi:2024flw , the curvature–induced deviation in a neutrino’s path—quantified by the deflection angle $\delta$—is intrinsically linked to the impact parameter $b$. Source $S$ produces neutrinos, which are then gravitationally lensed by the bumblebee black hole and finally detected at D. In the $(x,y)$ coordinate system, the physical distances from the source to the lens and the lens to the detector are denoted by the variables $r_{S}$ and $r_{D}$, respectively. In Ref 14 , the equation for the gravitational lensing of the bumblebee black hole has been derived carefully and strictly. This relationship is expressed through the following equation 14 :
@@
$=b^{2}\left(\dfrac{x_{D}}{r_{S}}+1-\dfrac{\pi\ell y_{D}}{2r_{S}}\right)-\dfrac{4Mby_{D}}{r_{S}}.$ (56)
We computed the ratio of $b_{1,2}^{2}/r_{S,D}^{2}$ by doing numerical calculations for $b$ in Eq. (VI), yielding the values displayed in Tab. 2. Four distinct angles $\phi=0,0.001,0.002,0.003$ were chosen as samples. We set the length to $\mathbf{fm}$ to correspond to the unit $\mathbf{GeV}$. The value of $b_{1,2}^{2}$ is far lower than that of $r_{S,D}^{2}$, as Tab. 2 makes evident. The correctness of employing the expansion $\mathcal{O}\left(b^{2}/r_{S,D}^{2}\right)$ in Eq. (40) is further demonstrated by the numerical solution of deflection angle Eq. (55).
Table 2: The ratio about $b_{1,2}^{2}/r_{S,D}^{2}$ with $r_{D}=10^{26}\,\mathrm{fm}$, $r_{S}=10^{31}\,\mathrm{fm}$ and $\ell=1\times 10^{-10}$.
$\phi=0$ $\phi=0.0015$ $\phi=0.003$
$b_{1}^{2}/r_{D}^{2}$ $5.99994\times 10^{-8}$ $2.36843\times 10^{-6}$ $9.11939\times 10^{-6}$
$b_{1}^{2}/r_{S}^{2}$ $5.99994\times 10^{-18}$ $2.36843\times 10^{-16}$ $9.11939\times 10^{-16}$
$b_{2}^{2}/r_{D}^{2}$ $5.99994\times 10^{-8}$ $1.51996\times 10^{-6}$ $3.94751\times 10^{-6}$
$b_{2}^{2}/r_{S}^{2}$ $5.99994\times 10^{-18}$ $1.51996\times 10^{-16}$ $3.94751\times 10^{-16}$
We analyzed how neutrino oscillations are affected by the spacetime of a Bumblebee black hole, focusing on gravitational corrections introduced through the Lorentz–violating parameter $\ell$, as we mentioned earlier. Assuming a circular trajectory of the detector around the Sun, where $x_{D}=r_{D}\cos\varphi$ and $y_{D}=r_{D}\sin\varphi$, we may numerically solve the quartic polynomial presented in Eq. (VI) inside the equatorial plane, yielding two positive real roots, $b_{1}$ and $b_{2}$, for each $\varphi$. Each term in Eq. (VI) is delineated by Eqs. (42)-(47), with conclusions derived by replacing the specified values of $m_{1}$, $m_{2}$, $b_{1}$ and $b_{2}$. Oscillation patterns were compared between the Bumblebee scenario ($\ell\neq 0$) and the Schwarzschild case ($\ell=0$), with Figs. 4, 5, and 6 highlighting how the azimuthal angle $\varphi$ and the mass–squared difference $\Delta m^{2}$ influence flavor transition probabilities.
@@
The influence of $\ell$ is detailed in Fig. 5. Increasing $\ell$ from $1\times 10^{-10}$ to $3\times 10^{-10}$ slightly changes the oscillation frequency but noticeably increases the whole probability of conversion. This aspect results from a promotion in the effective phase due to radial stretching of the spacetime. It is worthy to be highlight that, when the Kalb–Ramond fields takes into account instead, it did not produce such radial rescaling, showing no comparable suppression Shi:2025rfq .
On the other hand, Fig. 6 addresses the role of the absolute mass scale. As the lightest neutrino mass $m_{1}$ increases from 0 to $0.02\,\mathrm{eV}$, the amplitude decreases for the normal hierarchy and increases for the inverted one—behavior also seen in Kalb–Ramond gravity. However, the two models differ at $\varphi=0$ due to distinct modifications of the metric: Bumblebee spacetime preserves $\mathfrak{g}_{tt}$, whereas Kalb–Ramond modifies it via a $1/(1-\ell)$ factor, changing local energy measurements and initial oscillation conditions.
@@
Furthermore, as shown in the previous sections of this paper, the Lorentz–violating vector field of bumblebee gravity alters only the radial metric component $g_{rr}$, effectively stretching path lengths and increasing the oscillation phase; this leads to $\sim 20\%$ changes in the $\nu_{e}\!\to\!\nu_{\mu}$ conversion amplitude relative to the Schwarzschild case and a hierarchy–dependent modulation that reaches $\gtrsim 12\%$ across the azimuthal range. Such differences become observationally relevant in environments where neutrinos graze a compact object: (i) a core–collapse supernova whose proto–neutron star promptly forms a black-hole remnant, (ii) the hyper–accreting disk left after a binary–neutron‐star merger–both bright MeV sources that Hyper–Kamiokande or DUNE would record with $\mathcal{O}(10^{5})$ events at 10 kpc, yielding sub-percent statistical errors, and (iii) TeV–PeV neutrinos emitted along blazar jets, whose trajectories skirt the central super-massive black hole and are detectable by IceCube–Upgrade. Because the predicted flavor ratios differ from standard oscillations by $\sim 10\!-\!20\%$, a net experimental accuracy of $\lesssim 5\%$ in flavor identification suffices for a $3\sigma$ test; this benchmark lies within the projected capabilities of JUNO, DUNE and Hyper–K for MeV energies and IceCube–Gen2 for the high-energy channel.
@@ VII Conclusion
Propagation effects were governed by a reduced oscillation phase $\tilde{\Phi}_{k}=\frac{m_{k}^{2}}{2E_{0}}\left(1+\frac{\ell}{2}\right)(r_{D}+r_{S})\left(1-\frac{b^{2}}{2r_{D}r_{S}}+\frac{2M}{r_{D}+r_{S}}\right)$, so higher values of $\ell$ lengthened the oscillation scale along both radial and lensed paths. Numerical simulations revealed that raising $\ell$ from $1\times 10^{-10}$ to $3\times 10^{-10}$ left the azimuthal period almost unchanged but suppressed the $\nu_{e}\to\nu_{\mu}$ transition amplitude by up to twenty percent, an effect traced to the radial stretching of the optical metric. We pointed out that the increment of the oscillation amplitude was the relevant observable. For neutrinos coming from distant astrophysical sources, the rapid oscillations were averaged out, making the phase unobservable. Still, the increased amplitude in the flavor conversion probability, produced by Lorentz violation, remained as a stable and potentially measurable effect.
More so, the inverted mass ordering consistently produced higher conversion probabilities than the normal one with changing the lightest mass from 0 to $0.02\,\mathrm{eV}$. For phase calculations, we employed Schwarzschild-like coordinates for coordinate dependency, which hold true at the astrophysical scales utilized ($r_{S}=10^{5}r_{D}\gg 2M$). Near-event-horizon effects ($r\to 2M$) were not included in these parameter selections. The phase definition itself was geometrically invariant and independent of particular coordinate choices, as Ref. neu54 thoroughly demonstrated. The authors highlighted that the metric’s adjustments to actual distance and energy were what caused changes in the oscillation phase, not the fictitious idea of a ”gravitational phase”. We think that extreme conditions, including neutron star merging events, can detect the oscillation alterations suggested in Figs. 4-6. Our numerical simulations have investigated the region of $\ell\geq 10^{-10}$, where IceCube-Gen2’s sensitivity may be able to detect Lorentz symmetry breaking signals.
@@
In the future, more realistic techniques will need to incorporate quantum field theory, even though the toy model presented in this study can be considered a first step. Moreover, gravitational spin-flip transitions of neutrinos and gravitational decoherence effects must be taken into account in order to completely comprehend how gravitational influences affect neutrino oscillations. Although the consequences of quantum decoherence may be overlooked in the weak lensing limit, as noted in Ref. neu53 , it is a constraint for neutrino long-range transmission. Detectors need to be able to regulate wave packets with great accuracy when the lensing effect produces mass-dependent oscillation modes.